A sealed rectangular box measuring 6x15x17 cm contains 510 sugar cubes, each measuring 1x1x1 cm. The number of cubes touching he box is
Answer by Alan3354(69443)

A sealed rectangular box measuring 6x15x17 cm contains 510 sugar cubes, each measuring 1x1x1 cm. The number of cubes touching the box is
=================
If the box is full:
Total cubes = 6*15*17 = 1530
Find how many are 1 cube away from the faces.
That's a box that's 4x13x15 = 780.
1530 - 780 = 750
----------------------
With only 510 cubes in the box, several arrangements are possible.